Four objects are projected from a spring gun. Their masses and velocities are given in the following table: 7. Velocity (m/s) 10 4 Object Mass (kg) 2 4 8 2 Rank these objects by their momentum from greatest to least. A light-weight bowling ball has a mass of about 3 kg. A standard ping-pong ball has a mass of 0.003 kg. Assume that these balls have the same momentum, p 0.3 kg*m/s. a. Which ball requires a greater Impulse to stop? 8. b. Which ball has the greater velocity?
Answer #1

NOTE:

Momentum is defined as  

p = mv

Where, m is the mass and v is the velocity of the object. And

Impulse is defined as

  Impulse = Change in momentum

(7) From the given table:

Momentum of object A is

  PA-1 10-10 kg.m/s x

Momentum of object B is

P = 2 × 4 = 8 kg·m/s

Momentum of object C is

  Pe = 4× 3 = 12 kg. m/s

Momentum of object D is

PD = 8 × 2 = 16 kg·m/s

Hence, P_{D} > P_{C}> P_{A}> P_{B}

(8) According to the question, both the balls have same momentum i.e. p = 0.3 kg.m/s .

(a) If both the balls are going to stop then the change in momentum in both the case will be same (as both balls have same momentum). And we know that :

  Impulse = Change in momentum = 0.3 kg·m/s (for both cases)

Hence both ball will require same impulse to stop.

(b) Given,  

Mass of bowling ball = 3 kg

Mass of ping - pong ball = 0.003 kg

Momentum of both the balls = 0.3 kg.m/s

Now, we know that

Momentum = Mass x Velocity

or, Velocity = Momentum / Mass

So, Velocity of bowling ball = 0.3 / 3 = 0.1 m/s

And, velocity of ping - pong ball = 0.3 / 0.003 = 100 m/s

Hence, the ping - pong ball has greater velocity.

For any doubt please comment and please give an up vote. Thank you.